University Car Wash built a deluxe car wash across the street from campus. The new machines cost $246,000 including installation. The company estimates that the equipment will have a residual value of $27,000. University Car Wash also estimates it will use the machine for six years or about 12,000 total hours. Actual use per year was as follows:

Year Hours Used
1 2,800
2 1,900
3 2,000
4 2,000
5 1,800
6 1,500

2. Prepare a depreciation schedule for six years using the double-declining-balance method.
Answer #1

Answer-2)-

Depreciation Schedule-Double Declining Balance Method
YEAR COST $ Depreciation Rate Depreciation $ Accumulated Depreciation $ Book Value $
1 246000 33.33% 81992 81992 164008
2 33.33% 54664 136656 109344
3 33.33% 36444 173100 72900
4 33.33% 24298 197398 48602
5 33.33% 16199 213597 32403
6 33.33% 5403 219000 27000

Explanation- Double Declining balance depreciation is calculated using the following formula:

Depreciation = Depreciation Rate * Book Value of Asset

Depreciation rate is given by the following formula:

Depreciation Rate = Accelerator *Straight Line Rate

Straight-line Depreciation Rate = 1/6 = 0.1667 = 16.67%
Declining Balance Rate = 2*16.67% = 33.33%